# Model-free High Dimensional Mediator Selection with False Discovery Rate Control

### v1

@@ 5 Discussions
The ADNI data that support the findings in this paper can be accessed through the National Institute of Mental Health Data Archive (https://adni.loni.usc.edu/data-samples/adni-data/). The R package is available at https://github.com/RunqiuWang22/Generalized_Knockoff_Mediator_Selection.
@@
    We can run over a range of $\lambda$ values decreasing from +$\infty$+ ∞ (a fully sparse model) to 00 (a fully dense model) and define $Z_{j}$ as the maximum $\lambda$ such that $\widehat{\beta}_{j}(\lambda)\neq 0$. If there is no $\lambda$ such that $\widehat{\bm{\beta}}_{j}(\lambda)\neq 0$, then we will simply define $Z_{j}$ as 0.
@@
  • Absolute coefficient (glmnet): We can use $|\widehat{\beta}_{j}(\lambda)|$ from the penalized generalized linear regression or the penalized Cox regression model of $Y$ on $[\mathbf{X}\ \widetilde{\mathbf{X}}]$ with either a specific $\lambda$ value or a $\widehat{\lambda}$ estimated from cross-validation.
  • Standardized coefficient (glmnet): We can also use the standardized regression coefficients $|\widehat{\beta}_{j}(\lambda)|/\widehat{\text{SE}}(\widehat{\beta}_{j}(\lambda))$ from the penalized generalized linear regression or the penalized Cox regression model of $Y$ on $[\mathbf{X}\ \widetilde{\mathbf{X}}]$ with either a specific $\lambda$ value or a $\widehat{\lambda}$ estimated from cross-validation.
  • Order of selection (glmnet): We can use the minimum $\lambda$ such that the regression coefficient becomes 0 from the penalized generalized linear regression or the penalized Cox regression model of $Y$ on $[\mathbf{X}\ \widetilde{\mathbf{X}}]$, or the reciprocal of the order of each variable to be included in the model when increasing the number of variables to be selected.
  • Variable importance factor: We can use the variable importance factors from the random forest fitting of $Y$ on $[\mathbf{X}\ \widetilde{\mathbf{X}}]$ with either fixed tuning parameters or tuning parameters selected from cross-validation.
@@
Let $\mathbf{W}=f([\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}],[\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}])$ where $f$ is an OSFF. If $[Z^{a}_{j},\widetilde{Z}^{a}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{a}_{j},Z^{a}_{j}]$ for all $j\in\mathcal{H}_{a}$, $[Z^{b}_{j},\widetilde{Z}^{b}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{b}_{j},Z^{b}_{j}]$ for all $j\in\mathcal{H}_{b}$, and $[\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}]$ and $[\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}]$ are independent to each other, then for all $j\in\mathcal{H}$ and $\epsilon_{j}\sim\{\pm 1\}$ for all $j\in\mathcal{H}$. Then $W_{j}\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}W_{j}\cdot\epsilon_{j}$.
@@
For any $j\in\mathcal{H}$, if $j\in\mathcal{H}_{a}$, then we have $[Z^{a}_{j},\widetilde{Z}^{a}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{a}_{j},Z^{a}_{j}]$. Since $f$ is OSFF, we have $-W_{j}=f([\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}]_{Swap\{j\}},[\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}])_{j}$. Since $[\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}]$ and $[\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}]$ are independent, and $[Z^{a}_{j},\widetilde{Z}^{a}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{a}_{j},Z^{a}_{j}]$, we have
@@
Together, we have $W_{j}\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}-W_{j}$, and thus we have $W_{j}\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}W_{j}\cdot\epsilon_{j}$. Otherwise, if $j\in\mathcal{H}_{b}$, then we have $[Z^{b}_{j},\widetilde{Z}^{b}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{b}_{j},Z^{b}_{j}]$. Given $f$ is OSFF, we have $-W_{j}=f([\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}],[\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}]_{Swap\{j\}})_{j}$. Since $[\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}]$ and $[\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}]$ are independent, and $[Z^{b}_{j},\widetilde{Z}^{b}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{b}_{j},Z^{b}_{j}]$, we have
@@
The first part of the proof follows exactly the same as Barber & Candès (2015). The filtration $\mathcal{F}_{k}$ contains the information of whether $k$ is null and non-null process is known exactly. If $k$ is non-null, then $M(k-1)=M(k)$ and if $k$ is null, we have
@@
$\mathbb{E}\left[{M(k-1)|\mathcal{F}_{k}}\right]$ $=$ $\frac{1}{V^{+}(k)+V^{-}(k)}\left[V^{+}(k)\frac{V^{+}(k)-1}{V^{-}(k)+1}+V^{-}(k)\frac{V^{+}(k)}{V^{-}(k)\vee 1}\right]$
@@
$=$ $\mathbb{E}\left[{\frac{\#\{j\leq\widehat{k}_{+}:p_{j}\leq 1/2,j\in\mathcal{H}\}}{1+\#\{j\leq\widehat{k}_{+}:p_{j}>1/2,j\in\mathcal{H}\}}\left(\frac{1+\#\{j\leq\widehat{k}_{+}:p_{j}>1/2,j\in\mathcal{H}\}}{\#\{j\leq\widehat{k}_{+}:p_{j}\leq 1/2\}\vee 1}\right)\mathbbm{1}_{\widehat{k}_{+}>0}}\right]$
$\leq$ $\mathbb{E}\left[{\frac{\#\{j\leq\widehat{k}_{+}:p_{j}\leq 1/2,j\in\mathcal{H}\}}{1+\#\{j\leq\widehat{k}_{+}:p_{j}>1/2,j\in\mathcal{H}\}}}\right]q\leq q,$
@@
$\mathbb{E}\left[{\frac{V}{R+q^{-1}}}\right]$
$=$ $\mathbb{E}\left[{\frac{\#\{j\leq\widehat{k}_{0}:p_{j}\leq 1/2,j\in\mathcal{H}\}}{1+\#\{j\leq\widehat{k}_{0}:p_{j}>1/2,j\in\mathcal{H}\}}\left(\frac{1+\#\{j\leq\widehat{k}_{0}:p_{j}>1/2,j\in\mathcal{H}\}}{\#\{j\leq\widehat{k}_{0}:p_{j}\leq 1/2\}+q^{-1}}\right)\mathbbm{1}_{\widehat{k}_{0}>0}}\right]$
$\leq$ $\mathbb{E}\left[{\frac{\#\{j\leq\widehat{k}_{0}:p_{j}\leq 1/2,j\in\mathcal{H}\}}{1+\#\{j\leq\widehat{k}_{0}:p_{j}>1/2,j\in\mathcal{H}\}}}\right]q\leq q,$
@@
To prove Theorem 1, we first show that by the construction of the Model-X knockoff and the choice of compatible test statistics, we have $[Z^{a}_{j},\widetilde{Z}^{a}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{a}_{j},Z^{a}_{j}]$ for all $j\in\mathcal{H}_{a}$ and $[Z^{b}_{j},\widetilde{Z}^{b}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{b}_{j},Z^{b}_{j}]$ for all $j\in\mathcal{H}_{b}$. Since we can include $\mathbf{V}$ in the knockoff construction to make $[M_{j},\widetilde{M}_{j},\mathbf{V}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{M}_{j},M_{j},\mathbf{V}]$ for any $j\in\mathcal{H}_{a}$, we have the statistics from the marginal regression of $X$ on $[M_{j},\widetilde{M}_{j},\mathbf{V}]$ have the form of
$[Z^{a}_{j},\widetilde{Z}^{a}_{j}]=g([M_{j},\widetilde{M}_{j},\mathbf{V}]^{\top}[M_{j},\widetilde{M}_{j},\mathbf{V}],[M_{j},\widetilde{M}_{j},\mathbf{V}]^{\top}X).$
Since $[M_{j},\widetilde{M}_{j},\mathbf{V}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{M}_{j},M_{j},\mathbf{V}]$. For $j\in\mathcal{H}_{a}$, we have $[M_{j},\widetilde{M}_{j}]X=0$, so we have $Z^{a}_{j}\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}\widetilde{Z}^{a}_{j}$.
Also, since we have $[\mathbf{M},\widetilde{\mathbf{M}},\mathbf{V}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\mathbf{M},\widetilde{\mathbf{M}},\mathbf{V}]_{Swap(S)}$ for any $S\subset\mathcal{H}_{b}$, the Model-X knockoff results from Candès et al. (2018) and Dai et al. (2024) can be applied to obtain the exchangeability for $[Z^{b}_{j},\widetilde{Z}^{b}_{j}]$. When linear model is assumed, we can make a projection of $Y$, $X$, and $M_{j}$ on the orthogonal space of $\mathbf{V}$ for path-b. So the knockoff generated using residuals based on fixed knockoff construction and test statistics that compatible with fixed knockoff construction satisfy the exchangeability requirement for path-b by Barber & Candès (2015).
Since we use data splitting, we have $[\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}]$ and $[\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}]$ are independent. Define p-values $p_{j}=1$ if $W_{j}<0$ and $p_{j}=1/2$ if $W_{j}>0$, then Lemma B1 guarantee that we have $p_{j}\geq\text{Uniform}[0,1]$ and are independent from nonnulls. Since we assumed that the path-a statistics satisfy the weak dependence assumption, we have $p_{j}$s are weakly correlated. Applying Lemma B3, we obtain the mFDR and FDR control results. ∎

### v2

@@ 5 Discussions
Method Selected Mediator Effect (95% CI; p-value)
GKMS (RF)* Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; <<0.001)
$\beta$ = -0.157 (-0.388, -0.228; <<0.001))
Mediation = 0.063 (0.013, 0.092; <<0.001)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.213 (-0.458, -0.293;<<0.001)
Mediation =0.092 (0.051, 0.137; <<0.001)
Volume of Hippocampus (R) $\alpha$ = -0.551 (-0.658, -0.444; <<0.001)
$\beta$ = -0.162 (-0.393, -0.236; <<0.001)
Mediation = 0.026 (0.004, 0.072; 0.02)
Thick Avg of Angular Gyrus (L) $\alpha$ = 0.23 (0.115, 0.346; <<0.001)
$\beta$ = -0.135 (-0.342, -0.201; <<0.001)
Mediation = -0.024 (-0.045, -0.005; <<0.001)
ThickAvg of Sulcus (L) $\alpha$ = 0.192 (0.082, 0.302; <<0.001)
$\beta$ = -0.165 (-0.405, -0.234; <<0.001)
Mediation =-0.016 (-0.041, 0.003; 0.08)
GKMS (Lasso) Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; <<0.001)
$\beta$ = -0.132 (-0.193; -0.071;<<0.001)
Mediation = 0.069 (0.034, 0.104; <<0.001)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.178 (-0.238; -0.118;<<0.001)
Mediation = 0.09 (0.054, 0.126; <<0.001)
Volume of Amygdala (R) $\alpha$ = -0.596 (-0.706, -0.487; <<0.001)
$\beta$ = -0.111 (-0.171; -0.052;<<0.001)
Mediation = 0.066 (0.029, 0.104; 0.001)
Volume of Inf.Lat.Vent1 (R) $\alpha$ = -0.483(-0.591, -0.376; <<0.001)
$\beta$ = 0.15 (0.089; 0.211;<<0.001)
Mediation = -0.072 (-0.106, -0.039; <<0.001)
Volume of Brain Stem $\alpha$ = -0.955(-1.061, -0.848; <<0.001)
$\beta$ = -0.121 (0.06; 0.183;<<0.001)
Mediation = -0.116 (-0.176, -0.056; <<0.001)
MediationFDR Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.376 (-0.438; -0.314;<<0.001)
Mediation = 0.19 (0.139, 0.241; <<0.001)
HIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.282 (-0.343, -0.222; <<0.001)
Mediation = 0.143 (0.099, 0.186; <<0.001)
Volume of Inf.Lat.Vent (R) $\alpha$ = -0.483(-0.591, -0.376; <<0.001)
$\beta$ = 0.135 (0.074, 0.196; <<0.001)
Mediation = -0.065 (-0.098, -0.032; <<0.001)
ThickAvg of Parahip2 (R) $\alpha$ = 0.123 (0.007, 0.24; 0.037)
$\beta$ = -0.128 (-0.185, -0.072; <<0.001)
Mediation = -0.016 (-0.032, 0.001; 0.059)
eHIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.376 (-0.438; -0.314;<<0.001)
Mediation = 0.19 (0.139, 0.241; <<0.001)
  • *For GKMS (RF) method, $\beta$ effects are the importance scores with direction derived from partial dependence analysis comparing outcome predictions at the first and third quartiles of mediator values. Mediation effects are estimated using Monte Carlo integration as described by Imai et al. (2010) and confidence intervals are obtained via bootstrap.
  • Coefficient estimation is performed using structural equation modeling (SEM) via the lavaan R package (Rosseel, 2012).
  • 1 Inf.Lat.Vent refers to the Inferior Lateral Ventricle.
Method Selected Mediator Effect (95% CI; p-value)
GKMS (RF)* Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; <<0.001)
$\beta$ = -0.118 (-0.364, -0.186; <<0.001)
Mediation = 0.041 (0.01, 0.073; 0.016)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.171 (-0.411, -0.249; <<0.001)
Mediation = 0.102 (0.039, 0.127; <<0.001)
ThickAvg of Angular Gyrus(L) $\alpha$ = 0.233 (0.118, 0.349; <<0.001)
$\beta$ = -0.164 (-0.418, -0.222 <<0.001)
Mediation = -0.026 (-0.057, -0.005; 0.02)
$\beta$ = -0.177 (-0.422, -0.235; <<0.001)
Mediation =-0.004 (-0.032, 0.008; 0.22)
Volume of white matter in ITL 2(L) $\alpha$ = -0.895 (-1.004, -0.785; <<0.001)
$\beta$ = 0.006 (-0.155, 0.185; 0.9)
Mediation =-0.026 (-0.057, 0.022; 0.34)
GKMS (Lasso) Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; <<0.001)
$\beta$ = -0.083 (-0.145, -0.02; 0.009)
Mediation = 0.043 (0.009, 0.077; 0.012)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.139 (-0.201, -0.078; <<0.001)
Mediation = 0.07 (0.036, 0.105; <<0.001)
ThickAvg of Angular Gyrus(L) $\alpha$ = 0.233 (0.118, 0.349; <<0.001)
$\beta$ = -0.099 (-0.157, -0.041; 0.001)
Mediation = -0.023 (-0.041, -0.005; 0.011)
ThickAvg of S.temporal.sup(R) $\alpha$ = 0.092 (-0.023, 0.207; 0.117)
$\beta$ = -0.167 (-0.226, -0.109; <<0.001)
Mediation = -0.015 (-0.035, 0.005; 0.131)
Volume of Choroid Plexus (L) $\alpha$ = -0.719 (-0.826, -0.611; <<0.001)
$\beta$ = 0.081 (0.019, 0.143; 0.011)
Mediation = -0.058 (-0.104, -0.013; 0.012)
HIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.324 (-0.387, -0.261; <<0.001)
Mediation = 0.164 (0.116, 0.211; <<0.001)
eHIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = -0.324 (-0.387, -0.261; <<0.001)
Mediation = 0.164 (0.116, 0.211; <<0.001)
$\beta$ = -0.063 (-0.123, -0.004; 0.036)
Mediation = 0.013 (-0.001, 0.027; 0.078)
ThickAvg of Pole temporal(R) $\alpha$ = 0.259 (0.143, 0.376; <<0.001)
$\beta$ = -0.084 (-0.146, -0.023; 0.007)
Mediation = -0.022 (-0.041, -0.003; 0.022)
  • We use version ADAS-13 representing the results of ADAS-Cog.
  • *For GKMS (RF) method, $\beta$ effects are the importance scores with direction derived from partial dependence analysis comparing outcome predictions at the first and third quartiles of mediator values. Mediation effects are estimated using Monte Carlo integration as described by Imai et al. (2010) and confidence intervals are obtained via bootstrap.
  • Coefficient estimation is performed using structural equation modeling (SEM) via the lavaan R package (Rosseel, 2012).
Method Selected Mediator Effect (95% CI; p-value)
GKMS (RF)* Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; <<0.001)
$\beta$ = 0.129 (0.184, 0.375; <<0.001)
Mediation = -0.063 (-0.109, -0.016; 0.001)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ =0.165 (0.256, 0.409; <<0.001)
Mediation = -0.077 (-0.142, -0.026; <<0.001)
Volume of Amygdala (R) $\alpha$ = -0.596 (-0.706, -0.487; <<0.001)
$\beta$ = 0.157 (0.224, 0.396; <<0.001)
Mediation =-0.036 (-0.075, 0.001; 0.04)
Volume of Hippocampus (R) $\alpha$ = -0.551 (-0.658, -0.444; <<0.001)
$\beta$ = 0.132 (0.2, 0.366; <<0.001)
Mediation = -0.076 (-0.122, -0.029; 0.032)
ThickAvg of G-front-middle1 (L) $\alpha$ = 0.29 (0.175, 0.404; <<0.001)
$\beta$ = 0.136 (0.206, 0.392; <<0.001)
Mediation = 0.028 (0.013, 0.056; <<0.001)
GKMS (Lasso) Volume of Hippocampus (L) $\alpha$ = -0.52 (-0.627, -0.414; <<0.001)
$\beta$ = 0.111 (0.048, 0.174; 0.001)
Mediation = -0.058 (-0.093, -0.023; 0.001)
Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = 0.184 (0.122, 0.246; <<0.001)
Mediation = -0.093 (-0.13, -0.056; <<0.001)
Volume of Amygdala (R) $\alpha$ = -0.596 (-0.706, -0.487; <<0.001)
$\beta$ = 0.068 (0.007, 0.129; 0.028)
Mediation = -0.041 (-0.078, -0.004; 0.032)
Volume of Inf.Lat.Vent2 (R) $\alpha$ = -0.483(-0.591, -0.376; <<0.001)
$\beta$ = -0.088 (-0.15, -0.026; 0.006)
Mediation = 0.042 (0.011, 0.074; 0.008)
Volume of Medial Orbitofrontal (L) $\alpha$ = -0.702 (-0.816, -0.588; <<0.001)
$\beta$ = -0.147 (-0.206, -0.088; <<0.001)
Mediation = 0.103 (0.059, 0.148; <<0.001)
MediationFDR Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = 0.336 (0.273, 0.399; <<0.001)
Mediation = -0.17 (-0.218, -0.121; <<0.001)
HIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = 0.219 (0.158, 0.281; <<0.001)
Mediation = -0.111 (-0.15, -0.072; <<0.001)
ThickAvg of S.temporal.sup3 (L) $\alpha$ = 0.12 (0.007, 0.233; 0.037)
$\beta$ = 0.246 (0.187, 0.305; 0)
Mediation = 0.03 (0.001, 0.058; 0.043)
eHIMA Volume of Amygdala (L) $\alpha$ = -0.505 (-0.614, -0.397; <<0.001)
$\beta$ = 0.336 (0.273, 0.399; <<0.001)
Mediation = -0.17 (-0.218, -0.121; <<0.001)
  • *For GKMS (RF) method, $\beta$ effects are the importance scores with direction derived from partial dependence analysis comparing outcome predictions at the first and third quartiles of mediator values. Mediation effects are estimated using Monte Carlo integration as described by Imai et al. (2010) and confidence intervals are obtained via bootstrap.
  • Coefficient estimation is performed using structural equation modeling (SEM) via the lavaan R package (Rosseel, 2012).
  • 2 Inf.Lat.Vent refers to the Inferior Lateral Ventricle.
@@
The ADNI data that support the findings in this paper can be accessed through the National Institute of Mental Health Data Archive (https://adni.loni.usc.edu/data-samples/adni-data/). The R code is available at https://github.com/RunqiuWang22/Generalized_Knockoff_Mediator_Selection.
@@
    We can run over a range of $\lambda$ values decreasing from $\infty$ (a fully sparse model) to 0 (a fully dense model) and define $Z_{j}$ as the maximum $\lambda$ such that $\widehat{\beta}_{j}(\lambda)\neq 0$. If there is no $\lambda$ such that $\widehat{\bm{\beta}}_{j}(\lambda)\neq 0$, then we will simply define $Z_{j}$ as 0.
@@
  • Absolute coefficient (glmnet): We can use $|\widehat{\beta}_{j}(\lambda)|$ from the penalized generalized linear regression or the penalized Cox regression model of $Y$ on $\mathbf{X}\ \widetilde{\mathbf{X}}$] with either a specific $\lambda$ value or a $\widehat{\lambda}$ estimated from cross-validation.
  • Standardized coefficient (glmnet): We can also use the standardized regression coefficients $|\widehat{\beta}_{j}(\lambda)|/\widehat{\text{SE}}(\widehat{\beta}_{j}(\lambda))$ from the penalized generalized linear regression or the penalized Cox regression model of $Y$ on $\mathbf{X}\ \widetilde{\mathbf{X}}$] with either a specific $\lambda$ value or a $\widehat{\lambda}$ estimated from cross-validation.
  • Order of selection (glmnet): We can use the minimum $\lambda$ such that the regression coefficient becomes 0 from the penalized generalized linear regression or the penalized Cox regression model of $Y$ on $\mathbf{X}\ \widetilde{\mathbf{X}}$], or the reciprocal of the order of each variable to be included in the model when increasing the number of variables to be selected.
  • Variable importance factor: We can use the variable importance factors from the random forest fitting of $Y$ on $\mathbf{X}\ \widetilde{\mathbf{X}}$] with either fixed tuning parameters or tuning parameters selected from cross-validation.
@@
Let $\mathbf{W}=f([\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}],[\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}])$ where $f$ is an OSFF. If $Z^{a}_{j},\widetilde{Z}^{a}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{a}_{j},Z^{a}_{j}]$ for all $j\in\mathcal{H}_{a}$, $Z^{b}_{j},\widetilde{Z}^{b}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{b}_{j},Z^{b}_{j}]$ for all $j\in\mathcal{H}_{b}$, and $\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}]$ and $\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}]$ are independent to each other, then for all $j\in\mathcal{H}$ and $\epsilon_{j}\sim\{\pm 1\}$ for all $j\in\mathcal{H}$. Then $W_{j}\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}W_{j}\cdot\epsilon_{j}$.
@@
For any $j\in\mathcal{H}$, if $j\in\mathcal{H}_{a}$, then we have $Z^{a}_{j},\widetilde{Z}^{a}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{a}_{j},Z^{a}_{j}]$. Since $f$ is OSFF, we have $W_{j}=f([\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}]_{Swap\{j\}},[\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}])_{j}$. Since $\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}]$ and $\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}]$ are independent, and $Z^{a}_{j},\widetilde{Z}^{a}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{a}_{j},Z^{a}_{j}]$, we have
@@
Together, we have $W_{j}\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}-W_{j}$, and thus we have $W_{j}\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}W_{j}\cdot\epsilon_{j}$. Otherwise, if $j\in\mathcal{H}_{b}$, then we have $Z^{b}_{j},\widetilde{Z}^{b}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{b}_{j},Z^{b}_{j}]$. Given $f$ is OSFF, we have $W_{j}=f([\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}],[\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}]_{Swap\{j\}})_{j}$. Since $\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}]$ and $\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}]$ are independent, and $Z^{b}_{j},\widetilde{Z}^{b}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{b}_{j},Z^{b}_{j}]$, we have
@@
The first part of the proof follows exactly the same as Barber & Candès (2015). The filtration $\mathcal{F}_{k}$ contains the information of whether $k$ is null and non-null process is known exactly. If $k$ is non-null, then M(k1)=M(k)M(k-1)=M(k) and if $k$ is null, we have
@@
$\mathbb{E}\left[{M(k-1)|\mathcal{F}_{k}}\right]$
$=$ $\frac{1}{V^{+}(k)+V^{-}(k)}\left[V^{+}(k)\frac{V^{+}(k)-1}{V^{-}(k)+1}+V^{-}(k)\frac{V^{+}(k)}{V^{-}(k)\vee 1}\right]$
@@
$=\mathbb{E}\left[\frac{\#\left\{j\leq\widehat{k}_{+}:p_{j}\leq\frac{1}{2},\ j\in\mathcal{H}\right\}}{1+\#\left\{j\leq\widehat{k}_{+}:p_{j}>\frac{1}{2},\ j\in\mathcal{H}\right\}}\right$.
$\quad\left.\times\left(\frac{1+\#\left\{j\leq\widehat{k}_{+}:p_{j}>\frac{1}{2},\ j\in\mathcal{H}\right\}}{\#\left\{j\leq\widehat{k}_{+}:p_{j}\leq\frac{1}{2}\right\}\vee 1}\right)\mathbbm{1}_{\widehat{k}_{+}>0}\right]$
$\leq\mathbb{E}\left[\frac{\#\left\{j\leq\widehat{k}_{+}:p_{j}\leq\frac{1}{2},\ j\in\mathcal{H}\right\}}{1+\#\left\{j\leq\widehat{k}_{+}:p_{j}>\frac{1}{2},\ j\in\mathcal{H}\right\}}\right]q\leq q.$
@@
$\mathbb{E}\left[\frac{V}{R+q^{-1}}\right]$ $=\mathbb{E}\left[\frac{\#\left\{j\leq\widehat{k}_{0}:p_{j}\leq\frac{1}{2},\ j\in\mathcal{H}\right\}}{1+\#\left\{j\leq\widehat{k}_{0}:p_{j}>\frac{1}{2},\ j\in\mathcal{H}\right\}}\right$.
$\quad\left.\times\left(\frac{1+\#\left\{j\leq\widehat{k}_{0}:p_{j}>\frac{1}{2},\ j\in\mathcal{H}\right\}}{\#\left\{j\leq\widehat{k}_{0}:p_{j}\leq\frac{1}{2}\right\}+q^{-1}}\right)\mathbbm{1}_{\widehat{k}_{0}>0}\right]$
$\leq\mathbb{E}\left[\frac{\#\left\{j\leq\widehat{k}_{0}:p_{j}\leq\frac{1}{2},\ j\in\mathcal{H}\right\}}{1+\#\left\{j\leq\widehat{k}_{0}:p_{j}>\frac{1}{2},\ j\in\mathcal{H}\right\}}\right]q\leq q.$
@@
To prove Theorem 1, we first show that by the construction of the Model-X knockoff and the choice of compatible test statistics, we have $Z^{a}_{j},\widetilde{Z}^{a}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{a}_{j},Z^{a}_{j}]$ for all $j\in\mathcal{H}_{a}$ and $Z^{b}_{j},\widetilde{Z}^{b}_{j}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{Z}^{b}_{j},Z^{b}_{j}]$ for all $j\in\mathcal{H}_{b}$. Since we can include $\mathbf{V}$ in the knockoff construction to make $M_{j},\widetilde{M}_{j},\mathbf{V}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{M}_{j},M_{j},\mathbf{V}]$ for any $j\in\mathcal{H}_{a}$, we have the statistics from the marginal regression of $X$ on $M_{j},\widetilde{M}_{j},\mathbf{V}]$ have the form of
$Z^{a}_{j},\widetilde{Z}^{a}_{j}]=g([M_{j},\widetilde{M}_{j},\mathbf{V}]^{\top}[M_{j},\widetilde{M}_{j},\mathbf{V}],[M_{j},\widetilde{M}_{j},\mathbf{V}]^{\top}X).$
Since $M_{j},\widetilde{M}_{j},\mathbf{V}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\widetilde{M}_{j},M_{j},\mathbf{V}]$. For $j\in\mathcal{H}_{a}$, we have $M_{j},\widetilde{M}_{j}]X=0$, so we have $Z^{a}_{j}\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}\widetilde{Z}^{a}_{j}$.
Also, since we have $\mathbf{M},\widetilde{\mathbf{M}},\mathbf{V}]\stackrel{{\scriptstyle\mathclap{\mbox{d}}}}{{=}}[\mathbf{M},\widetilde{\mathbf{M}},\mathbf{V}]_{Swap(S)}$ for any $S\subset\mathcal{H}_{b}$, the Model-X knockoff results from Candès et al. (2018) and Dai et al. (2024) can be applied to obtain the exchangeability for $Z^{b}_{j},\widetilde{Z}^{b}_{j}]$. When linear model is assumed, we can make a projection of $Y$, $X$, and $M_{j}$ on the orthogonal space of $\mathbf{V}$ for path-b. So the knockoff generated using residuals based on fixed knockoff construction and test statistics that compatible with fixed knockoff construction satisfy the exchangeability requirement for path-b by Barber & Candès (2015).
Since we use data splitting, we have $\mathbf{Z}^{a},\widetilde{\mathbf{Z}}^{a}]$ and $\mathbf{Z}^{b},\widetilde{\mathbf{Z}}^{b}]$ are independent. Define p-values $p_{j}=1$ if $W_{j}<0$ and $p_{j}=1/2$ if $W_{j}>0$, then Lemma B1 guarantee that we have $p_{j}\geq\text{Uniform}[0,1]$ and are independent from nonnulls. Since we assumed that the path-a statistics satisfy the weak dependence assumption, we have $p_{j}$s are weakly correlated. Applying Lemma B3, we obtain the mFDR and FDR control results. ∎